The Brisbane Manufacturing Company produces a single model of a CD player. Each player is sold for $187 with a resulting contribution margin of $80. Brisbane's management is considering a change in its quality control system. Currently, Brisbane spends $42,000 a year to inspect the CD players. An average of 2,100 units turn out to be defective - 1,680 of them are detected in the inspection process and are repaired for $75. If a defective CD player is not identified in the inspection process, the customer who receives it is given a full refund of the purchase price. The proposed quality control system involves the purchase of an x-ray machine for $190,000. The machine would last for four years and would have salvage value at that time of $21,000. Brisbane would also spend $430,000 immediately to train workers to better detect and repair defective units. Annual inspection costs would increase by $21,000. This new control system would reduce the number of defective units to 370 per year. 305 of these defective units would be detected and repaired at a cost of $44 per unit. Customers who still received defective players would be given a refund equal to 120% of the purchase price. Questions 1 & 2 [0 points; unlimited tries] 1. What is the Year 2 cash flow if Brisbane keeps using its current system? $-246,540 You are correct. Your receipt no. is 156-9049 Help: Receipt Previous Tries 2. What is the Year 2 cash flow if Brisbane replaces its current system? $-91,006 You are correct. Your receipt no. is 156-4100 Help: Receipt Previous Tries Questions 3 & 4 [5 points each; 5 tries each] [NOTE: When computing present values, use the present value factors from the tables on page 118 in the Coursepack] 3. Assuming a discount rate of 6%, what is the net present value if Brisbane keeps using its current system? $-854,261 You are correct. Your receipt no. is 156-770 Help: Receipt Previous Tries 4. Assuming a discount rate of 6%, what is the net present value if Brisbane replaces its current system?
Answer #1

cash flow for current system

Total defective units-detected

=2100-1680

=420

year 2 cash flow

Annual inspection cost -repairing cost *units repaired- refunded units

= -42000 -(75$*1680)-420*187

=-42000-126000-78540

= -$246,540

new system cash flow in year 2

annual inspection cost -repairing cost*units repaired - refunded units

= -[42000+21000increase]-305*44-[370-305]*187*120%

= -63000-13420-14586

= $-910,06

NPV is difference between present value of cash inflow and cash outflow

positive net present value means that the project is better

CURRENT SYSTEM

YEAR cash flow pv factor at 6% present value
1-4 -$246,540 3.465[1/1.06]1+[1/1.06]2+[1/1.06]3+[1/1.06]4 -$854,261

new system

YEAR cash flow pv factor at 6% presnt value
1-4 $-91,006 3.465 -$315,336
0 $-190,000 1 -$190,000
0 -430000 1 -$430,000
4 21000 0.792 $16,632
NET PRESENT VALUE -$918,704